# Spectral Dataset of Young Type Ib Supernovae and their Time-evolution

### v1

@@ 4.7 SN 2019odp: transitioning from SN Ic-bl to SN Ib
SN 2019odp is an interesting transitional object, which shows features like SNe Ic-bl before the V-band maximum light (t 0) but transitions to Type Ib afterwards (see also Schweyer et al. 2023 with their own set of spectra for this SN). SN 2019odp was discovered (Nordin et al., 2019) by ZTF on UT 2019-08-21 09:19:53 (MJD 58716.39) with a non-detection 7 days prior and was initially classified as SNe Ic-bl on 23 August 2019 (MJD = 58718.2) by Brennan et al. (2019) as part of the ePESSTO+ survey (Smartt et al., 2015). Using presumably the same spectrum, Schweyer et al. (2023) found that the spectrum was observed 16 days before the g-band peak (t = -16d) and that it matches SNe Ic-bl. However, they reclassified SN 2019odp as Type Ib based on the spectrum taken 26 days after the g-band peak (t = 26d). Thus, Schweyer et al. (2024) updated the classification of SN 2019odp as Type Ib on TNS. We ran the early spectra and the spectrum closest to the V-band peak date through SNID. SNID found SN 2007bi, a hydrogen-poor super-luminous SN Ic (SLSNe Ic; Liu et al. 2017), as the best match for the earliest spectrum (t = -15.7d). For the subsequent three early spectra (t = -11.6d, -8.7d, -5.6d) , SNID found the best match with SN 1999bw (SNe Ic-bl; Patat et al. 2001). Liu et al. (2017) found that the average absorption velocities are similar for SLSNe Ic and SNe Ic-bl, suggesting high velocity features in the earliest spectrum of SN 2019odp. The best match for our spectrum at -2.7d is a normal SN Ic, SN 2013ge. For the spectrum closest to the V-band maximum light, taken on UT 2019-09-10 13:47:02 (t = 1.4d), SNID found that the best-fitting type is normal Ib, with 55.9 of the templates with rlap rlapmin corresponding to this type. Thus, we re-classify SN 2019odp as a normal SN Ib based on spectra very close to maximum light (at t = 1.4d) – 3 weeks earlier than Schweyer et al. (2023) did – and indicate its transitional nature in Tables 1 and 2 with ”Ic-bl -¿ Ib”. This transition from a SN Ic-bl to a SN Ib has been observed once before, in the famous SN 2008D, whose spectra resembled that of Type Ic-bl shortly after explosion, i.e., 15–10 days before the date of maximum light, but developed narrow absorption lines of helium by the date of maximum light (Soderberg et al. 2008; Mazzali et al. 2008, Modjaz et al. 2009), similar to SN 2019odp.
@@ 4.8 SN 2016bau
SN 2016bau was discovered on UT 2016-03-13 23:22:33 (MJD 57460.97) with a non-detection 9 days prior (Arbour, 2016). It was classified as an SN Ib on TNS using a classification spectrum obtained one day after discovery (Benetti et al., 2016). We ran our spectrum closest to the V-band maximum, taken on UT 2016-03-29 10:25:44 (t = -1.7d) through SNID. SNID found that the best-fitting type is normal SN Ib, with 48.1 of the templates with rlap rlapmin corresponding to this type. For the earliest spectrum taken on UT 2016-03-15 05:24:59 (t = -15.9d), SNID found that the best-fitting type is normal SN Ib, with 62.5% of the templates with rlap rlapmin corresponding to this type. We note that Aryan et al. (2021) published their own extensive data set on this SN (including eight optical spectra, two of which were taken before maximum light), with a date of maximum calculation that is consistent with ours (see Section 3).
@@ 5 Constructing SNID Templates from our Sample of Young SNe Ib, Including Published Spectra
In order to improve the classification by SNID, it is crucial to enrich the SNID library by adding more classified SN templates of early SNe Ib. We created SNID templates using the 82 photospheric spectra of the 8 SNe Ib published in this work. In addition, we included the 19 spectra of SN 2019odp published in Schweyer et al. (2023) and 6 spectra of SN 2016bau published in Aryan et al. (2021) when creating the SNID templates of those SNe. We used the logwave777https://people.lam.fr/blondin.stephane/software/snid/howto.html#logwave function in SNID to create spectral templates that are compatible with SNID 5.0. Our work adds 38 new SNe Ib spectra with ¡ 0 to the SNID library, increasing the early SNe Ib spectra by 43. Including previously published spectra for SN 2019odp and SN 2016bau as mentioned above, the total increase is 54. SNID templates for the SNe Ib presented in this paper are publicly available via the METAL GitHub repository888https://github.com/metal-sn/SESNtemple/tree/master, along with the previously published Modjaz Group sample (MGS) SNID templates, including those of Young SN Ic from Williamson et al. (2023).
@@ 6 Conclusions
In the era of transient surveys with extensive sky coverage and rapid cadences such as ATLAS and ZTF, and with LSST on the horizon, prompt and accurate classification of young SNe holds significant importance for subsequent follow-up observations. In this paper, we present a data set of 82 photospheric spectra of 8 SNe Ib, which includes a total of 38 spectra observed before the V-band maximum. This spectral data set is publicly available via WISeREP. This data set increases the number of published SNe Ib with at least three premaximum spectra by 60 and the number of available premaximum SNe Ib spectra in the SNID template library by 43.
@@
1) As mentioned above, almost half of our sample had a different type than listed in TNS or an evolving type over time, similar to the sample of Young SNe Ic we published in Williamson et al. (2023). Thus, we suggest that the community keep observing and classifying SNe Ib over time when they are discovered young - and in general young CCSNe, including Young SNe Ic - so that the community can be alerted about any SNe that change types or develop peculiarities. In addition, since one three of our SN sample had time-evolving types (SN 2019odp, SN 2021ukt, and 2022nyo), with additional ones in the literature (e.g., SN 2008D, SN 2022crv) we suggest that alternative SN identification methods are used that can capture the SN’s amount of transitioning between different SESN subclasses - one such method was suggested by us in Williamson et al. (2019b), where we developed a PCA method for the photospheric spectra of SESNe which can naturally capture transition SNe (as applied to SN 2022crv, see Figure 17 in Dong et al. 2023b).
2) Incorporating early SN spectra into classification libraries as well as training sets for machine learning tools can improve the classification of young SNe. We suggest that the spectra of young SNe Ib, such as those presented here, are included in the spectral libraries of the various SN identification codes, e.g. SNID (Blondin et al., 2006), Superfit (Howell et al., 2005) and GELATO (Harutyunyan et al., 2008). For the benefit of the community, we produce and make public the SNID templates of our Young SNe Ib spectral dataset in this paper.
Our release of these young spectra, which offer information about the outermost layers of SNe Ib ejecta, will hopefully facilitate work on constraining the characteristics and mass-loss mechanism of the SN progenitor and on investigating different spectral characteristics, such as measuring line velocities and strength, to quantify diversity in SNe Ib at early times.
@@ Acknowledgements.
Appendix A Observation Log

### v2

@@ 4.7 SN 2019odp: Transitioning from SN Ic-bl to SN Ib
SN 2019odp is an interesting transitional object, which shows features like SNe Ic-bl before the V-band maximum light (t 0) but transitions to type Ib afterwards (see also Schweyer et al. 2023 with their own set of spectra for this SN). SN 2019odp was discovered (Nordin et al., 2019) by ZTF on UT 2019-08-21 09:19:53 (MJD 58716.39) with a non-detection 7 days prior and was initially classified as SNe Ic-bl on UT 2019-08-23 05:23:25 (MJD = 58718.2) by Brennan et al. (2019) as part of the ePESSTO+ survey (Smartt et al., 2015). Using presumably the same spectrum, Schweyer et al. (2023) found that the spectrum was observed 16 days before the g-band peak (t = 16d) and that it matches SNe Ic-bl. However, they reclassified SN 2019odp as type Ib based on the spectrum taken 26 days after the g-band peak (t = 26d). Thus, Schweyer et al. (2024) updated the classification of SN 2019odp as type Ib on TNS. We ran the early spectra and the spectrum closest to the V-band peak date through SNID. SNID found SN 2007bi, a hydrogen-poor super-luminous SN Ic (SLSNe Ic; Liu et al. 2017), as the best match for the earliest spectrum (t = 15.7d). For the subsequent three early spectra (t = 11.6d, 8.7d, 5.6d) , SNID found the best match with SN 1999bw (SNe Ic-bl; Patat et al. 2001). Liu et al. (2017) found that the average absorption velocities are similar for SLSNe Ic and SNe Ic-bl, suggesting high velocity features in the earliest spectrum of SN 2019odp. The best match for our spectrum at 2.7d is a normal SN Ic, SN 2013ge. For the spectrum closest to the V-band maximum light, taken on UT 2019-09-10 13:47:02 (t = 1.4d), SNID found that the best-fitting type is normal Ib, with 55.9 of the templates with rlap rlapmin corresponding to this type. Thus, we reclassify SN 2019odp as a normal SN Ib based on spectra very close to maximum light (at t = 1.4d) – three weeks earlier than Schweyer et al. (2023) did – and indicate its transitional nature in Tables 1 and 3 with ”Ic-bl -¿ Ib”. This transition from a SN Ic-bl to a SN Ib has been observed once before, in the famous SN 2008D, whose spectra resembled that of type Ic-bl shortly after explosion, that is, 15–10 days before the date of maximum light, but developed narrow absorption lines of helium by the date of maximum light (Soderberg et al. 2008; Mazzali et al. 2008; Modjaz et al. 2009), similar to SN 2019odp.
@@ 4.8 SN 2016bau
SN 2016bau was discovered on UT 2016-03-13 23:22:33 (MJD 57460.97) with a non-detection 9 days prior (Arbour, 2016). It was classified as an SN Ib on TNS using a classification spectrum obtained one day after discovery (Benetti et al., 2016). We ran our spectrum closest to the V-band maximum, taken on UT 2016-03-29 10:25:44 (t = 1.7d) through SNID. SNID found that the best-fitting type is normal SN Ib, with 48.1 of the templates with rlap rlapmin corresponding to this type. For the earliest spectrum taken on UT 2016-03-15 05:24:59 (t = 15.9d), SNID found that the best-fitting type is normal SN Ib, with 62.5% of the templates with rlap rlapmin corresponding to this type. We note that Aryan et al. (2021) published their own extensive dataset on this SN (including eight optical spectra, two of which were taken before maximum light), with a date of maximum calculation that is consistent with ours (see Section 3).
@@ 5 Constructing SNID templates from our sample of young SNe Ib, including published spectra
In order to improve the classification by SNID, it is crucial to enrich the SNID library by adding more classified SN templates of young SNe Ib. We created SNID templates using the 82 photospheric spectra of the eight SNe Ib published in this work. In addition, we included the 19 spectra of SN 2019odp published in Schweyer et al. (2023) and 6 spectra of SN 2016bau published in Aryan et al. (2021) when creating the SNID templates of those SNe. We note that all spectra in our sample are labeled as type Ib in our SNID templates, with the exception of type-changing SNe, such as SN 2022nyo, SN 2021ukt, and SN 2019odp, which are labeled as Ib-pec (peculiar type Ib). These objects are classified as peculiar type Ib due to their type-changing behavior. We also emphasize that the label ”Ib-pec” in the SNID library can refer to various peculiarities (peculiar light curve behavior, changing type over time, etc.) shown by a SN. We encourage readers to explore the details of these specific SNe to better understand the unique characteristics that lead to their classification as peculiar.
We used the logwave888https://people.lam.fr/blondin.stephane/software/snid/howto.html#logwave function in SNID to create spectral templates that are compatible with SNID 5.0. Our work adds 38 new SNe Ib spectra with ¡ 0 to the SNID library, increasing the early SNe Ib spectra by 43. Including previously published spectra for SN 2019odp and SN 2016bau as mentioned above, the total increase is 54. SNID templates for the SNe Ib presented in this paper are publicly available via the METAL GitHub repository999https://github.com/metal-sn/SESNtemple/tree/master, along with the previously published Modjaz Group sample (MGS) SNID templates, including those of young SNe Ic from Williamson et al. (2023).
@@ 6 Conclusions
In the era of transient surveys with extensive sky coverage and rapid cadences such as ATLAS and ZTF, and with LSST on the horizon, the prompt and accurate classification of young SNe holds significant importance for subsequent follow-up observations. In this paper, we present a dataset of 82 photospheric spectra of eight SNe Ib, which includes a total of 38 spectra observed before the V-band maximum. This spectral dataset is publicly available via WISeREP. This dataset increases the number of published SNe Ib with at least three pre-maximum spectra by 60 and the number of available pre-maximum SNe Ib spectra in the SNID template library by 43.
@@
1) As mentioned above, almost half of our sample had a different type than listed in TNS or an evolving type over time, similar to the sample of young SNe Ic we published in Williamson et al. (2023). Thus, we suggest that the community keep observing and classifying SNe Ib over time when they are discovered young - and in general young CCSNe, including young SNe Ic - so that the community can be alerted about any SNe that change types or develop peculiarities. In addition, since three of the eight SNe our sample had time-evolving types (SN 2019odp, SN 2021ukt, and 2022nyo), with additional ones in the literature (e.g., SN 2008D and SN 2022crv) we suggest that alternative SN identification methods be used that can capture the SN’s amount of transitioning between different SESN subclasses - one such method was suggested by us in Williamson et al. (2019), where we developed a PCA method for the photospheric spectra of SESNe that can naturally capture transition SNe (as applied to SN 2022crv, see Figure 17 in Dong et al. 2024).
2) Incorporating early SN spectra into classification libraries as well as training sets for machine learning tools can improve the classification of young SNe. We suggest that the spectra of young SNe Ib, such as those presented here, be included in the spectral libraries of the various SN identification codes, such as SNID, Superfit, and GELATO. For the benefit of the community, we produce and make public the SNID templates of our young SNe Ib spectral dataset in this paper and release them via our METAL GitHub repository, formerly known as the SNYU github repository.
Our release of these young spectra, which offer information about the outermost layers of SNe Ib ejecta, will hopefully facilitate work on constraining the characteristics and mass-loss mechanism of the SN progenitor. Additionally, these spectra will enable investigating different spectral characteristics, such as measuring line velocities and strength, to quantify diversity in SNe Ib at early times.
@@ Acknowledgements.
Appendix A Spectral time-series plots
Appendix B Observation log